OpenStudy (anonymous):

PLEASE HELP!! Using the properties of exponents, simplify the expression. 5^1/2*5^1/2

OpenStudy (anonymous):

go to this website

OpenStudy (anonymous):

http://www.wolframalpha.com/ here

OpenStudy (anonymous):

just type the question and it will give the answer not always if it does not work try different ways

OpenStudy (mathstudent55):

\(a^m \cdot a^n = a^{m + n} \) To multiply two powers with the same base, write the same base and add the exponents.

OpenStudy (anonymous):

5^-2/2? What next?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Wait. It's not that. Is your problem \(\large 5^\frac{1}{2} \cdot 5^\frac{1}{2} \) ?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

If that is the problem, then all you need to do is add the exponents. What is 1/2 + 1/2 = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

No the problem is 5^1/2 * 5^-2/2 @mathstudent55

OpenStudy (mathstudent55):

That's not what you wrote above in your original post.

OpenStudy (mathstudent55):

Ok, let's do this version of the problem. Once again, remember, all you need to do is add exponents since the bases are the same.

OpenStudy (anonymous):

Yes I wrote the wrong thing on the original post. Sorry!

OpenStudy (mathstudent55):